# Counting spaces of functions on separable compact lines

### v1

@@ Abstract.
We investigate the following general problem, closely related to the problem of isomorphic classification of Banach spaces $C(K)$ of continuous real-valued functions on a compact space $K$, equipped with the standard supremum norm:
Let $\mathcal{K}$ be a class of compact spaces. How many isomorphism types of Banach spaces $C(K)$ of real-valued continuous functions on $K$ are there, for $K\in\mathcal{K}$?
@@
We show that, for the class $\mathcal{L}_{\omega_{1}}$ of separable compact linearly ordered spaces of weight $\omega_{1}$, the answer to the above question depends on additional set-theoretic axioms. In particular, assuming the continuum hypothesis, there are $2^{2^{\omega}}$ isomorphism types of $C(L)$, for $L\in\mathcal{L}_{\omega_{1}}$, and assuming a certain axiom proposed by Baumgartner, there is only one type.
@@
Primary 46B03, 46E15, 54F05, Secondary 54G12
The second and the third authors were partially supported by the NCN (National Science Centre, Poland) research grant no. 2020/37/B/ST1/02613.
@@ 1. Introduction
One of the most fundamental and natural challenges of mathematics is the problem of classifying objects within a given class. An important part of such a challenge is an attempt to determine how many isomorphism types of objects we have from that class.
In this paper, we consider the following case of such a general problem:
Let $\mathcal{K}$ be a class of compact spaces. How many isomorphism types of Banach spaces $C(K)$ of real-valued continuous functions on $K$, equipped with the standard supremum norm, do we have for $K\in\mathcal{K}$?
Recall that for an infinite compact space $K$, the weight of $K$ is equal to the weight and the density of $C(K)$. Hence, if the Banach spaces $C(K)$ and $C(L)$ are isomorphic, then the compact spaces $K$ and $L$ must have equal weights. It is also known that compact spaces $K$ and $L$ with isomorphic function spaces have equal cardinalities. Therefore, it is natural that in the problem of counting isomorphism types of spaces $C(K)$, we restrict ourselves to the classes of compacta $K$ of fixed weight or fixed cardinality.
The classical results of Miljutin, Bessaga and Pełczyński give us a complete isomorphic classification of spaces $C(K)$ for metrizable compacta $K$, i.e., for compacta of countable weight. In particular, Bessaga and Pełczyński (cf. [32]) proved that there are exactly $\omega_{1}$ isomorphism types of $C(K)$, for countable (hence metrizable) $K$, and Miljutin (cf. [32]) showed that, for uncountable metrizable $K$, there is only one isomorphic type of $C(K)$ space. Summarizing, we have $\omega_{1}$ isomorphism types of spaces $C(K)$ for compacta $K$ of weight $\omega$ (regardless of the size of the continuum $2^{\omega}$). Recall here that we have $2^{\omega}$ topological types of compacta of weight $\omega$.
For the classes of nonmetrizable compacta, the chances of obtaining such complete classification results seem to be very slim, for a simple reason - we usually have too many isomorphism types of spaces $C(K)$. It is known that even if we restrict ourselves to some concrete classes of compact spaces of weight $2^{\omega}$, we still have $2^{2^{\omega}}$ isomorphism types of $C(K)$. In particular, it is true for the class of separable scattered compacta of scattered height 3 (cf. [6]), or the class of separable, compact linearly ordered spaces (cf. [21]).
Note that, for compacta $K$ of weight $\kappa$, the cardinality of isomorphism types of $C(K)$ is bounded by $2^{\kappa}$ (cf. remarks in the last section).
In this paper, we concentrate on the case of compact spaces of weight $\omega_{1}$. We prove that we have $2^{\omega_{1}}$ isomorphism types of $C(K)$, for $K$ of such weight. Actually, we prove the following more general result:
@@
Most of our investigations deal with the class of separable, compact linearly ordered spaces (in short: separable compact lines). Banach spaces $C(K)$ associated with such compacta $K$ play an important role in Banach space theory, they provided many interesting examples (see [15],[16], [17], [27]), and were an object of investigations in numerous papers (see [2], [7], [14], [20], [21], [23], [9]).
Let $\mathcal{L}_{\kappa}$ be the class of separable, compact linearly ordered spaces of weight $\kappa$. We show that the cardinality of isomorphism types of $C(L)$, for $L\in\mathcal{L}_{\omega_{1}}$, depends on additional set-theoretic axioms. In particular, assuming the continuum hypothesis, there are $2^{2^{\omega}}$ isomorphism types of $C(L)$, for $L\in\mathcal{L}_{\omega_{1}}$. More precisely, we prove the following:
@@
Let us recall that the first example of a class $\mathcal{K}$ of compact spaces of weight $\omega_{1}$, such that the cardinality of isomorphism types of $C(K)$, for $K\in\mathcal{K}$, depends on additional set-theoretic axioms, was given in [6] ($\mathcal{K}$ is the class of separable scattered compacta of scattered height 3).
@@ 2. Preliminaries
2.1. Set-theoretical terminology
We denote by $cf(\alpha)$ the cofinality of the ordinal $\alpha$. Let $\kappa$ be a regular cardinal, i.e., $cf(\kappa)=\kappa$. Recall that a set $A\subseteq\kappa$ is a club set if it is unbounded in $\kappa$ and closed in the order topology. A set $S\subseteq\kappa$ is stationary if it has a nonempty intersection with all club subsets of $\kappa$. By $\omega_{1}$ we denote the smallest uncountable cardinal and usually write $2^{\omega}$ for the continuum, the cardinality of the set of real numbers. In general, our set-theoretical terminology follows Jech’s book [13].
@@ 2.2. Banach spaces of continuous functions and their duals
The dual space $C(K)^{*}$, due to the Riesz representation theorem, can be seen as $M(K)$, the space of signed Radon measures on $K$ with absolute variation as the norm. There is a similar representation of $C_{0}(X)^{*}$ for any locally compact space $X$, where every element of $C_{0}(X)^{*}$ can be represented as integration with respect to a signed Radon measure on the one-point compactification $\alpha X$ of $X$ that vanishes at the added point (see Theorem 18.4.1 of [32]).
In the space $M(K)$, equipped with the weak* topology, there is a linearly dense topological copy of $K$, which is the set of Dirac delta measures $\Delta_{K}=\{\delta_{k}:k\in K\}$ (we will sometimes identify $K$ and $\Delta_{K}$). This, in particular, means that if the space $K$ is separable, then $M(K)$ is also separable in the weak* topology. We write $supp(\mu)$ for the support of the measure $\mu\in M(K)$.
@@
The following fact is standard, see e.g. [32, Proposition 21.5.8]. Note that all hyperplanes of a Banach space are mutually isomorphic, see [11, Exercise 2.9].
@@ 2.3. Scattered spaces
For a scattered space $X$, by Cantor-Bendixson height $ht(X)$ of $X$ we mean the minimal ordinal $\alpha$ such that the Cantor-Bendixson derivative $X^{(\alpha)}$ of the space $X$ is empty. The Cantor-Bendixson height of a compact scattered space is always a nonlimit ordinal.
Let us make a general remark about the bounded linear functionals on $C_{0}(X)$ for any locally compact scattered space $X$. By the Rudin theorem (Theorem 19.7.6 of [32]), if $\alpha X$ is scattered, then all Radon measures on the one-point compactification $\alpha X$ are purely atomic. From the Riesz representation theorem mentioned above follows that any bounded linear functional $\varphi$ on $C_{0}(X)$ can be seen as
@@
where $\sum_{n\in\omega}|a_{n}|<\infty$ and $\{x_{n}:n\in\omega\}\subseteq X$ are distinct.
@@ 2.4. Separable compact linearly ordered spaces
and consider this set with the order topology given by the lexicographic order. If $A=(0,1)$ and $K=I$, then the space $\mathbb{K}=K_{A}$ is the well-known double arrow space. Some authors use this name for the space $I_{I}$; others call the space $I_{I}$ the split interval. For any sets $A,K$ as above, the space $K_{A}$ is a separable compact line of topological weight $|A|$ (if the set $A$ is infinite), and thus it is nonmetrizable whenever the set $A$ is uncountable. If $A$ is dense in $K$, then the space $K_{A}$ is zero-dimensional. It turns out that spaces of the form $K_{A}$ exhaust the class of separable compact lines:
@@
Due to Ostaszewski’s result, in this paper we can assume that any separable compact line is of the form $K_{A}$.
3. Counting isomorphism types of $C(K)$ of given weight
@@
The following space, known, at least for $\kappa=\omega_{1}$, as the ladder system space, seems to belong to the topological folklore since the early 1970s. Perhaps its most known early appearances are in [8] and [31].
@@
Consider an uncountable regular cardinal $\kappa$ and $S\subseteq E^{\kappa}_{\omega}$. Let $\mathcal{L}=\{L_{\alpha}:\alpha\in S\}$ be a system of ladders, i.e., for $\alpha\in S$ we demand
@@
We define the ladder system space $X_{S}(\mathcal{L})=X_{S}$ on $\kappa$:
@@
  • The basic neighbourhoods of $\alpha\in S$ are declared as all sets of the form
@@
By the Stone-Weierstrass theorem, $C_{0}(X_{S})$ is a hyperplane of $C(KL_{S})$. The space $KL_{S}$ is scattered, so it admits a nontrivial convergent sequence. The rest follows from Fact 2.1. ∎
@@
Suppose that $\kappa$ is an uncountable regular cardinal and that $R,S\subseteq E_{\omega}(\kappa)$ are two sets such that $S\setminus R$ is stationary in $\kappa$. Then there is no bounded linear operator from $C_{0}(X_{R})$ into $C_{0}(X_{S})$ with a dense range.
@@
Suppose that $T:C_{0}(X_{R})\rightarrow C_{0}(X_{S})$ is a bounded linear operator. We will show that it does not have a dense range. Consider the adjoint operator $T^{*}:C_{0}(X_{S})^{*}\rightarrow C_{0}(X_{R})^{*}$. Let $F,G:\kappa\rightarrow\kappa$ be such that
@@
Indeed, when $f\in C_{0}(X_{R})$, then for every $\varepsilon>0$ there is $\beta<\alpha$ such that all values of $f|[\beta,\alpha)$ are in the interval $(-\varepsilon,\varepsilon)$. Otherwise, there would be a strictly increasing sequence $(\alpha_{n})_{n\in\omega}$ of points in $\alpha$ satisfying $|f(\alpha_{n})|\geq\varepsilon$. As the point $\alpha$ is isolated in $X_{R}$ and the basic neighbourhood of other points in $X_{R}$ have bounded intersections with $\alpha$, the sequence $(\alpha_{n})_{n\in\omega}$ is divergent to infinity, which is a contradiction.
It follows from the above and the fact that $T^{*}(\delta_{\beta_{n}})$ form a bounded sequence of measures that the sequence $T^{*}(\delta_{\beta_{n}})(f)$ converge to 0 for each $f\in C_{0}(X_{R})$. Thus, we have that the sequence $(T^{*}(\delta_{\beta_{n}}))_{n\in\omega}$ weakly converges to 0 in $C_{0}(X_{S})^{*}$ and $T^{*}(\delta_{\alpha})=0$ as needed to conclude that $T$ does not have dense range. ∎
@@
Fix $\kappa$ as in the statement. By a theorem of Solovay (Theorem 8.10 of [13] also cf. Lemma 8.8 of [13]) there is a pairwise disjoint family $\{S_{\xi}:\xi<\kappa\}$ of stationary subsets of $E_{\omega}^{\kappa}$. For every $A\subset\kappa$ consider
@@ 4. Topological properties of separable compact lines
First, we will show that for any cardinal number $\kappa,\omega\leq\kappa\leq 2^{\omega}$, there are $2^{\kappa}$ pairwise nonhomeomorphic separable compact lines $K_{A}$ of weight $\kappa$ (recall that, for infinite $A$, the weight of $K_{A}$ is equal to $|A|$).
We start with the following observation of van Douwen [35, Section 10].
@@
The above proposition together with a standard fact that each subset of $I$ can be homeomorphic to at most continuum many subsets of $I$ easy implies that for a fixed separable compact line $K_{A}$ we have at most $2^{\omega}$ many separable compact lines $L_{B}$ homeomorphic to $K_{A}$. This immediately allows us to obtain the following variation of van Douwen’s result that there are $2^{2^{\omega}}$ separable compact lines.
@@
Enumerate the family of all components of $I\setminus K$ as $\{(a_{n},b_{n}):n\in N\}$, where $N=\omega$ or $N\in\omega$. Observe that, for each $n\in N$, the set $\{a_{n},b_{n}\}$ is disjoint from $A$, since $K_{A}$ has no isolated points.
Now consider an equivalence relation on $K$ defined by $a_{n}\!\sim b_{n}$ for all $n\in N$ and let $q:K\to K/\!\!\sim$ be the quotient map. Notice that the relation $\sim$ is closed, as every increasing or decreasing sequence of $a_{n}$’s will have the same limit as a sequence of $b_{n}$’s with the same indexes. One can easily check that the resulting quotient space $K/\!\!\sim$ is a linearly ordered connected metrizable compact space, hence it is homeomorphic to the unit interval (cf. [10, 6.3.2]). Denote by $h:K/\!\!\sim\ \to I$ the homeomorphism satisfying $h(q(0))=0$. One can verify that $K_{A}$ can be identified with $I_{B}$ where $B=h[q[A\cup\{a_{n}:n\in N\}]]$. ∎
@@
for $s\in 2^{<\omega},b\in\{0,1\}$. Since $h$ is order-preserving, $Q$ is dense in $I$, and $\{q_{s}^{0},q_{s}^{1}:s\in 2^{<\omega}\}$ is dense in $C$, we can extend $h$ to an order isomorphism $\widetilde{h}:I_{A}\to C_{B}$, where $B=h\big[A\setminus Q\big]$. ∎

### v3

@@ Abstract.
We investigate the following general problem, closely related to the problem of isomorphic classification of Banach spaces $C(K)$ of continuous real-valued functions on a compact space $K$, equipped with the supremum norm:
Let $\mathcal{K}$ be a class of compact spaces. How many isomorphism types of Banach spaces $C(K)$ are there, for $K\in\mathcal{K}$?
@@
We show that, for the class $\mathcal{L}_{\omega_{1}}$ of separable compact linearly ordered spaces of weight $\omega_{1}$, the answer to the above question depends on additional set-theoretic axioms. In particular, assuming the continuum hypothesis, there are $2^{\omega_{1}}$ isomorphism types of $C(L)$, for $L\in\mathcal{L}_{\omega_{1}}$, and assuming a certain axiom proposed by Baumgartner, there is only one type.
@@
Primary 03E35, 03E65, 03E75, 46B03, 46E15, 54F05
The second-named and the third-named authors were partially supported by the NCN (National Science Centre, Poland) research grant no. 2020/37/B/ST1/02613.
@@ 1. Introduction
One of the most fundamental and natural challenges of mathematics is the problem of classifying objects within a given class. An important part of such a challenge is an attempt to determine how many isomorphism types of objects we have in that class.
In this paper, we consider the following case of this general problem:
Let $\mathcal{K}$ be a class of compact spaces. How many isomorphism types of Banach spaces $C(K)$ of real-valued continuous functions on $K$, equipped with the supremum norm, do we have for $K\in\mathcal{K}$?
Recall that for an infinite compact space $K$, the weight of $K$ is equal to the weight and the density of $C(K)$. Hence, if the Banach spaces $C(K)$ and $C(L)$ are isomorphic, then the compact spaces $K$ and $L$ must have equal weights. It is also known that compact spaces $K$ and $L$ with isomorphic function spaces have equal cardinalities [7]. Therefore, it is natural that in the problem of counting isomorphism types of spaces $C(K)$, we restrict ourselves to the classes of compacta $K$ of fixed weight or fixed cardinality.
The classical results of Miljutin, Bessaga and Pełczyński give us a complete isomorphic classification of spaces $C(K)$ for metrizable compacta $K$, i.e., for compacta of countable weight. In particular, Bessaga and Pełczyński (cf. [34]) proved that there are exactly $\omega_{1}$ isomorphism types of $C(K)$, for countable (hence metrizable) $K$, and Miljutin (cf. [34]) showed that, for uncountable metrizable $K$, there is only one isomorphic type of $C(K)$ space. Summarizing, we have $\omega_{1}$ isomorphism types of spaces $C(K)$ for compacta $K$ of weight $\omega$ (regardless of the size of the continuum $2^{\omega}$). Recall that we have $2^{\omega}$ topological types of compacta of weight $\omega$ [24].
For the classes of nonmetrizable compacta, the chances of obtaining such complete classification results seem to be very slim, for a simple reason - we usually have too many isomorphism types of spaces $C(K)$. It is known that even if we restrict ourselves to some concrete classes of compact spaces of weight $2^{\omega}$, we still have $2^{2^{\omega}}$ isomorphism types of $C(K)$. In particular, it is true for the class of separable scattered compacta of height 3 (cf. [6]), or the class of separable, compact linearly ordered spaces (cf. [23]).
Note that, for a compact $K$ of weight $\kappa$, the cardinality of isomorphism types of $C(K)$ is bounded by $2^{\kappa}$ (cf. remarks in the last section).
In this paper, we concentrate on the case of compact spaces of weight $\omega_{1}$. We prove that there are $2^{\omega_{1}}$ isomorphism types of $C(K)$, for $K$ of such weight. Actually, we prove the following more general result:
@@
Most of our investigations deal with the class of separable, compact linearly ordered spaces (in short: separable compact lines). Banach spaces $C(K)$ associated with such compacta $K$ play an important role in Banach space theory. They have provided many interesting examples (see [16], [17], [18], [29]), and have been an object of investigations in numerous papers (see [2], [8], [15], [21], [23], [25], [10]).
Let $\mathcal{L}_{\kappa}$ be the class of separable, compact linearly ordered spaces of weight $\kappa$ (for our topological terminology see [11]). We show that the cardinality of isomorphism types of $C(L)$, for $L\in\mathcal{L}_{\omega_{1}}$, depends on additional set-theoretic axioms. In particular, assuming the continuum hypothesis (i.e., $2^{\omega}=\omega_{1}$), there are $2^{2^{\omega}}$ isomorphism types of $C(L)$, for $L\in\mathcal{L}_{\omega_{1}}$. More precisely, we prove the following:
@@
The above corollary should be compared with a recent result of Kucharski [22] who proved that there exists a separable compact line $K$ of weight $2^{\omega}$ such that the Banach space $C(K)$ is not isomorphic to $C(K)\oplus C(K)$. In particular, this shows that Corollary 1.4 requires some additional set-theoretic hypothesis.
Let us recall that another example of a class $\mathcal{K}$ of compact spaces of weight $\omega_{1}$, such that the cardinality of isomorphism types of $C(K)$, for $K\in\mathcal{K}$, depends on additional set-theoretic axioms, was given in [6] ($\mathcal{K}$ is the class of separable scattered compacta of height 3).
@@ 2. Preliminaries
2.1. Set-theoretic terminology
We denote by $cf(\alpha)$ the cofinality of the ordinal $\alpha$. Let $\kappa$ be a regular cardinal, i.e., $cf(\kappa)=\kappa$. Recall that a set $A\subseteq\kappa$ is a club set if it is unbounded in $\kappa$ and closed in the order topology. A set $S\subseteq\kappa$ is stationary if it has a nonempty intersection with all club subsets of $\kappa$. By $\omega_{1}$ we denote the first uncountable cardinal and usually write $2^{\omega}$ for the continuum, the cardinality of the set of real numbers. In general, our set-theoretic terminology follows Jech’s book [14].
@@ 2.2. Banach spaces of continuous functions and their duals
The dual space $C(K)^{*}$, due to the Riesz representation theorem, can be seen as $M(K)$, the Banach space of signed Radon measures on $K$ with absolute variation as the norm. There is a similar representation of $C_{0}(X)^{*}$ for any locally compact space $X$, where every element of $C_{0}(X)^{*}$ can be represented as the integration with respect to a signed Radon measure on the one-point compactification $\alpha X$ of $X$ that vanishes at the added point (see Theorem 18.4.1 of [34]).
In the space $M(K)$, equipped with the weak* topology induced by $C(K)$, there is a linearly dense topological copy of $K$, which is the set of Dirac delta measures $\Delta_{K}=\{\delta_{x}:x\in K\}$ (we will sometimes identify $K$ and $\Delta_{K}$). This, in particular, means that if the space $K$ is separable, then $M(K)$ is also separable in the weak* topology. We write $supp(\mu)$ for the support of the measure $\mu\in M(K)$.
@@
The following fact is standard, see e.g. [34, Proposition 21.5.8]. Note that all hyperplanes of a Banach space are mutually isomorphic, see [12, Exercise 2.9].
@@
In general, our terminology concerning Banach spaces follows books [12] and [34].
@@ 2.3. Scattered spaces
For a scattered space $X$, by the Cantor-Bendixson height (or simply the height) $ht(X)$ of $X$ we mean the least ordinal $\alpha$ such that the Cantor-Bendixson derivative $X^{(\alpha)}$ of the space $X$ is empty. The height of a compact scattered space is always a nonlimit ordinal.
Let us make a general remark about the bounded linear functionals on $C_{0}(X)$ for any locally compact scattered space $X$. By the Rudin theorem (Theorem 19.7.6 of [34]), if $\alpha X$ is scattered, then all Radon measures on the one-point compactification $\alpha X$ are purely atomic. From the Riesz representation theorem mentioned above it follows that any bounded linear functional $\varphi$ on $C_{0}(X)$ is defined by
@@
where $\sum_{n\in\omega}|a_{n}|<\infty$ and $\{x_{n}:n\in\omega\}\subseteq X$ are distinct, and $f\in C_{0}(X)$.
@@ 2.4. Separable compact linearly ordered spaces
and consider this set with the order topology given by the lexicographic order on $0,1]\times\{0,1\}$. If $A=(0,1)$ and $K=I$, then the space $K_{A}$ is the well-known double arrow space. Some authors use this name for the space $I_{I}$; others call the space $I_{I}$ the split interval. For any sets $A,K$ as above, the space $K_{A}$ is a separable compact line of topological weight $|A|$ (if the set $A$ is infinite), and thus it is nonmetrizable whenever the set $A$ is uncountable. If $A$ is dense in $K$, then the space $K_{A}$ is zero-dimensional. It turns out that spaces of the form $K_{A}$ exhaust the class of separable compact lines:
@@
Due to Ostaszewski’s result, in this paper we may assume that any separable compact line is of the form $K_{A}$.
3. Counting isomorphism types of $C(K)$ of a given weight
@@
The following space, known, at least for $\kappa=\omega_{1}$, as the ladder system space, seems to belong to the topological folklore since the early 1970s. Perhaps its best-known early appearances are in [9] and [33].
@@
Consider an uncountable regular cardinal $\kappa$. Let $\mathcal{L}=\{L_{\alpha}:\alpha\in E^{\kappa}_{\omega}\}$ be a system of ladders, i.e., for $\alpha\in E^{\kappa}_{\omega}$ we demand
@@
For a subset $S$ of $E^{\kappa}_{\omega}$, we define the ladder system space $X_{S}(\mathcal{L})=X_{S}$ on $\kappa$:
@@
  • The basic neighborhoods of $\alpha\in S$ are declared as all sets of the form
@@
By the Stone-Weierstrass theorem, $C_{0}(X_{S})$ is a hyperplane of $C(KL_{S})$. The space $KL_{S}$ is scattered and infinite, so it admits a nontrivial convergent sequence. The rest follows from Fact 2.1. ∎
@@
Suppose that $\kappa$ is an uncountable regular cardinal and that $R,S\subseteq E_{\omega}^{\kappa}$ are two sets such that $S\setminus R$ is stationary in $\kappa$. Then there is no bounded linear operator from $C_{0}(X_{R})$ into $C_{0}(X_{S})$ with a dense range.
@@
Suppose that $T:C_{0}(X_{R})\rightarrow C_{0}(X_{S})$ is a bounded linear operator. We will show that it does not have a dense range. Consider the adjoint operator $T^{*}:C_{0}(X_{S})^{*}\rightarrow C_{0}(X_{R})^{*}$. Let $F,G:\kappa\rightarrow\kappa$ be nondecreasing functions such that
@@
Indeed, when $f\in C_{0}(X_{R})$, then for every $\varepsilon>0$ there is $\beta<\alpha$ such that all values of $f|[\beta,\alpha)$ are in the interval $(-\varepsilon,\varepsilon)$. Otherwise, there would be a strictly increasing and cofinal sequence $(\alpha_{n})_{n\in\omega}$ of points in $\alpha$ satisfying $|f(\alpha_{n})|\geq\varepsilon$. As the point $\alpha$ is isolated in $X_{R}$ and the basic neighborhood of other points in $X_{R}$ have bounded intersections with $\alpha$, the sequence $(\alpha_{n})_{n\in\omega}$ is divergent to infinity, which is a contradiction.
It follows from the above and the fact that $T^{*}(\delta_{\beta_{n}})$ form a bounded sequence of measures that the sequence $T^{*}(\delta_{\beta_{n}})(f)$ converges to 0 for each $f\in C_{0}(X_{R})$. Thus, we have that the sequence $(T^{*}(\delta_{\beta_{n}}))_{n\in\omega}$ weakly converges to 0 in $C_{0}(X_{S})^{*}$ and $T^{*}(\delta_{\alpha})=0$ as needed to conclude that $T$ does not have dense range. ∎
@@
Fix $\kappa$ as in the statement. By a theorem of Solovay (Theorem 8.10 of [14] also cf. Lemma 8.8 of [14]) there is a pairwise disjoint family $\{S_{\xi}:\xi<\kappa\}$ of stationary subsets of $E_{\omega}^{\kappa}$. For every $A\subseteq\kappa$ consider
@@ 4. Topological properties of separable compact lines
In this section we use our convention that we identify separable compact lines with the spaces of the form $K_{A}$, where $K\subseteq[0,1]$ and $A\subseteq K$, see subsection 2.4.
First, we will show that for any cardinal number $\kappa$ satisfying $\omega\leq\kappa\leq 2^{\omega}$, there are $2^{\kappa}$ pairwise nonhomeomorphic separable compact lines $K_{A}$ of weight $\kappa$ (recall that, for infinite $A$, the weight of $K_{A}$ is equal to $|A|$).
We start with the following observation of van Douwen [37, Section 10].
@@
The above proposition together with a standard fact that each subset of $I$ can be homeomorphic to at most continuum many subsets of $I$, and some routine cardinal calculations, easily implies that for a fixed separable compact line $K_{A}$ we have at most $2^{\omega}$ many separable compact lines $L_{B}$ homeomorphic to $K_{A}$ (cf. [37, Section 10], [23, Section 2]). This immediately allows us to obtain the following variation of van Douwen’s result that there are $2^{2^{\omega}}$ separable compact lines.
@@
Enumerate the family of all connectedness components of $I\setminus K$ as $\{(a_{n},b_{n}):n\in N\}$, where $N=\omega$ or $N\in\omega$. Observe that, for each $n\in N$, the set $\{a_{n},b_{n}\}$ is disjoint from $A$, since $K_{A}$ has no isolated points.
Now consider an equivalence relation on $K$ defined by $a_{n}\!\sim b_{n}$ for all $n\in N$ and let $q:K\to K/\!\!\sim$ be the quotient map. Notice that the relation $\sim$ is a closed relation (see [11]), as every increasing or decreasing sequence of $a_{n}$’s will have the same limit as a sequence of $b_{n}$’s with the same indicies. One can easily check that the resulting quotient space $K/\!\!\sim$ is a linearly ordered connected metrizable compact space, hence it is order isomorphic to the unit interval (cf. [11, 6.3.2]). Denote by $h:K/\!\!\sim\ \to I$ the order isomorphism satisfying $h(q(0))=0$. One can verify that $K_{A}$ can be identified with $I_{B}$ where $B=h[q[A\cup\{a_{n}:n\in N\}]]$. ∎
@@
for $s\in 2^{<\omega},b\in\{0,1\}$. Since $h$ is order-preserving (with respect to the lexicographic order on $Q\times\{0,1\}$), $Q$ is dense in $I$, and $\{q_{s}^{0},q_{s}^{1}:s\in 2^{<\omega}\}$ is dense in $C$, we can extend $h$ to an order isomorphism $\widetilde{h}:I_{A}\to C_{B}$, where $B\times\{1\}=\widetilde{h}\big[(A\setminus Q\big)\times\{1\}]$. ∎